OpenStudy (anonymous):

Find the coordinates of the midpoint between point A(-5, -9) and the origin. A. five-halves, nine-halves B. (-5, -9) C. negative five-halves, negative nine-halves D. (5, 9)

OpenStudy (anonymous):

to find the mid point we use the following formula \[x =(\frac{ x _{1} + x _{2} }{ 2 } )\] and for y \[y = \frac{ y _{1} + y _{2} }{ 2 }\]

OpenStudy (anonymous):

and since the co-ordinates of the origin are 0,0 => x = ( -5 + 0 ) / 2 = -5/2 and y = (-9 + 0 )/2 = -9/2 hence A is the correct option... :)

OpenStudy (anonymous):

Thanks!!!